the base of a triangle is 7cm greater than the height. the area is 22cm^2. find the height and length of the base.
Answer by nerdybill(7384)

the base of a triangle is 7cm greater than the height. the area is 22cm^2. find the height and length of the base.
.
Since
area of any triangle is (1/2)bh
where
b is base
h is height
.
Let h = height
then
h+7 = base
.
22 = (1/2)(h+7)h
44 = (h+7)h
44 = h^2 + 7h
0 = h^2 + 7h - 44
0 = (h+11)(h-4)
h = {-11, 4}
we can throw out the negative solution leaving:
h = 4 cm (height)
.
base:
h+7 = 4+7 = 11 cm